OpenStudy (anonymous):

How many molecules (or formula units) are in each sample? 74.41 g ( CCl4 )

OpenStudy (anonymous):

Just find number of moles= weight/Molecular weight. Then find moles * N where N is 6.022*10^23 or avogadro number. moles=74.41/(12+4*35.5)=0.483. Formula units= 0.483*6.022*10^23=3*10^23

OpenStudy (anonymous):

The answer is N=2.913*10^23 Thanks :)

OpenStudy (anonymous):

You're welcome :)